OpenStudy (anonymous):

How many atoms are in 2.56 mol of (NH4)3 PO4?

OpenStudy (anonymous):

Please quickly answer

OpenStudy (anonymous):

calculate the molar mass of the compound .!

OpenStudy (anonymous):

Multiply first by the Avogrados number.

OpenStudy (anonymous):

yes ,,,15.416 x 10^23 atoms

OpenStudy (anonymous):

but that is not the answer on the back. I get 1.54*10powerof24 but answer in the back says 3.08*10powerof25

OpenStudy (anonymous):

There are 2.56 x 6 x 10^23 molecules in a mole of the compound, but each molecule has a lot of atoms that need to be counted.

OpenStudy (anonymous):

explain more please

OpenStudy (anonymous):

Take the compound description and count how many atoms of each element make up one molecule of the compound. e.g., NH4 has 5 atoms

OpenStudy (anonymous):

\[2.56 \text{ mol} (NH_4)_3PO_4\times \frac{6.02 \times 10^{23} \text{ molecules}(NH_4)_3PO_4}{1 \text{ mol}(NH_4)_3PO_4} \times \frac{20 \text{ atoms}}{1 \text{ molecule}(NH_4)_3PO_4}\]

OpenStudy (anonymous):

ok thanks add me on snapchat johncenafan69

OpenStudy (anonymous):

how many atoms are in \((NH_4)_3 PO_4\)? well, \(3\times(1+4)+1+4=15+5=20\) atoms Avogadro's constant tells us that \(1\text{ mol}\) contains approximately \(6.022\times10^{23}\) molecules so the total number of atoms is just$$6.022\times10^{23}\cdot20=1.204\times10^{25}$$

Parth (parthkohli):

But what oldrin said is the number of atoms in one mole. Multiply that by 2.56 to get the answer.